OpenStudy (anonymous):

Find f^-1(x) given that f (x) is 2/3-4x

OpenStudy (anonymous):

f(x) = 2 x 1/3-4x let's break it down: 1. multiply 4 by x 2. subtract product from 3 3. divide 1 by the resut of step 2 4 and multiply resut of step 3 by 2 Now let's do the opposite: 1. divide 4 by x = 4/x 2. add product to 3= 3+4/x 3. multiply 1 by result = 3+4/x 4 divide resut of step 3 by 2 = 3+4x/2 Check: let x = 2 f(x) = 2/3 + 4(2) = 2/11 F^-1(x) = 3+4(2)/2 = 11/2

OpenStudy (anonymous):

\[ y=2/3-4x\\ 4 x = 2/3 -y\\ x= 1/6 -y/4\\ f^{-1}(x)=1/6 -x/4 \]